OpenStudy (anonymous):

what is the ratio of radioactive element to decay element after 2 half-lives?

thomaster (thomaster):

1/4 of the initial mass of the element 1x half life = 1/2 so 2 times the half live is 1/2 of 1/2 (1/2*1/2=1/4)